# Identifying Equivalent Training Dynamics

### v2

@@ 4.2 Identifying effect of width on fully connected neural network training
Computing the Koopman eigenvalues associated with training FCNs of varying width (Fig. 3D – see Appendix B for details), we find that narrow ($h=5$) and wide ($h=40$) FCNs have training dynamics that are non-conjugate, as their Koopman spectra are significantly different, with respect to the Wasserstein distance (Fig. 3E, F). This suggests that the training dynamics undergo a fundamental change as width increases. However, for FCNs with intermediate width ($h=10$), the training dynamics are more aligned with the wide FCNs (Fig. 3E, F), suggesting conjugate dynamical behavior. Thus, we conclude that the additional improvement in performance observed when increasing the network width from $h=10$ to $h=40$ comes more from an increase in capacity, than from a change in training dynamics. Identifying this was not possible by solely comparing the loss or weights (Fig. 3A–C), demonstrating the advantage of the Koopman-based approach for identifying equivalent and non-equivalent DNN training dynamics
@@ 4.3 Identifying dynamical transitions in convolutional neural network training
Prior work has argued that CNNs undergo transitions in their dynamics during the early part of training (i.e. the first several epochs), and that these transitions are similar across different CNN architectures [27]. However, dynamical systems based methods were not used for analysis. Instead, this observation relied on coarse-grained observables (e.g., training loss, magnitude of gradients) to define the transitions and to determine when they occur.
@@
To understand whether such results hold when considering the training dynamics at a finer-scale, we utilize our Koopman-based framework. To do this, we split the first epoch of training into windows of 100 training iterations. We compute the Koopman eigenvalues associated with dynamics that occur in each window and denote them by $\lambda_{T_{1}:T_{2}}$, where $T_{1}<T_{2}$ are the first and last training iteration in the window. We then measure the Wasserstein distance between all pairs of eigenvalues. This enables us to quantitatively assess transient dynamical behavior and identify when in the early phase of training the dynamics transition from one equivalence class to another. We apply our approach to LeNet [55], a simple CNN [^2] While “simple”, LeNet has over 200K parameters. trained on MNIST, and ResNet-20 [56], trained on CIFAR-10 (see Appendix C for details).
@@
We find that, for both LeNet and ResNet-20, the first 100 training iterations have the most distinct Koopman eigenvalues, as the Wasserstein distance between $\lambda_{0:100}$ and all other eigenvalues is large (Fig. 4A, B – bright yellow first column and row). In addition, for both LeNet and ResNet-20, the training dynamics become similar after $700$ training iterations, as the Wasserstein distance between $\lambda_{600:700}$ and $\lambda_{700:800}$ is small (Fig. 4A, B – dark blue square around diagonal in lower righthand corner). This is in agreement with the timeline found by Frankle et al. (2020) [27]. However, we additionally find that the dynamics that occur between 100 and 700 training iterations exhibit greater change for ResNet-20 than for LeNet, as there is a larger Wasserstein distance between Koopman eigenvalues. This suggests a difference in dynamical behavior between the architectures. By examining the Koopman eigenvalues associated with different training iteration windows, we find non-overlapping spectra (Fig. 4C). This suggests a lack of topological conjugacy between the training dynamics of ResNet-20 and LeNet, despite the fact that the general timeline in transitions in dynamics is similar between the architectures.
@@
Figure 4: Koopman-based framework enables identification of transitions in dynamics during the early phase of training for LeNet and ResNet-20. (A) Log10 Wasserstein distance between Koopman eigenvalues associated with LeNet training over windows of 100 training iterations during epoch 1. (B) Same as (A), but for ResNet-20 training. (C) Koopman eigenvalues associated with the dynamics that occur over training iterations 0–100, 400–500, and 600–700. Dashed line denotes unit circle.
@@
To understand how the training dynamics change over a larger span of training time, we perform the same analysis, but computing the Koopman eigenvalues from the dynamics that occur during each epoch. We find that, at this coarser grain scale, both architectures see a similar evolution of their training dynamics. In particular, we find that the first epoch has the most distinct dynamics (Fig. 5A, B – yellow first column and row), and the subsequent epochs have dynamics that become increasingly more similar (Fig. 5A, B – increasing size of dark blue blocks centered on the diagonal).
Figure 5: Koopman-based framework enables identification of transitions in dynamics across the training of LeNet and ResNet-20. (A)–(B) Same as Fig. 4(A)–(B), but for dynamics computed over individual epochs.
@@
Taken together, our Koopman-based analysis supports prior decomposition of the early phase of CNN training dynamics into regimes separated by transitions that occur at similar training iterations across architectures [27]. However, with a finer-scale resolution of the dynamics, we additionally find that LeNet and ResNet-20 have non-conjugate training, demonstrating that the exact dynamics are architecture specific.
@@ 4.4 Identifying non-conjugate training dynamics for Transformers that do and do not grok
Since the discovery that Transformers trained on algorithmic data (e.g. modular addition) undergo delayed generalization (or “grokking” – Fig. 6A) [28], considerable effort has been performed to understand how this arises. One particularly influential theory is that the norm of the weights at initialization plays a large role. In particular, it was shown that Transformers initialized at weights with a sufficiently small norm have training and test loss landscapes that are aligned, and Transformers initialized at weights with a sufficiently large norm have training and test loss landscapes that are mis-aligned [57]. Constraining the norm of the weights to be small prevents grokking, with train and test accuracy increasing at similar training iterations (Fig. 6B) [57].
Figure 6: Transformers that do and do not undergo grokking have early training dynamics that are conjugate. (A) Train and test loss, as a function of training steps, for a Transformer model that undergoes grokking. (B) Same as (A), but for a Transformer whose training is constrained to have a constant weight norm [57]. In this case, no grokking is observed. (C) In the first 100 training steps, little difference is seen between the test loss of Transformers with and without constrained training. Lines are mean and shaded area is $\pm$ standard deviation across 20 independently trained networks. (D) Koopman eigenvalues associated with the dynamics that occur over the first 100 training iterations for Transformers that do and not undergo grokking.
However, what role the training dynamics play in grokking remains less understood. In particular, the extent to which constraining the weight norm changes the training dynamics (which could shed additional light on grokking) has yet to be explored. We therefore compute the Koopman eigenvalues associated with the training of constrained and unconstrained Transformers on modular addition (Appendix D). We use the dynamics from the earliest part of training, namely the first 100 training iterations (Fig. 6C). We do this to avoid trivially seeing a difference given the small weight changes that Transformers which undergo grokking make when the training accuracy is high.
@@
These results suggest a non-conjugacy in the training dynamics of Transformers that do and do not undergo grokking. In particular, constraining the weight norm appears to lead to more stable training dynamics, which may be due to the selection of a better subnetwork to train [58]. Additionally, these results suggest that it might be possibly to identify grokking before it happens [59].
@@ 5 Discussion
Motivated by the need for rigorous methods that can determine when the training dynamics of DNN parameter trajectories are equivalent, we utilized advances in Koopman operator theory to develop framework that can identify topological conjugacies. This Koopman-based spectral identification of conjugate training dynamics is invariant to permutations of the ordering of DNN parameter labels (Eq. 5), a needed feature for methods used to study DNN training dynamics [15, 16, 17]. By applying our approach to the dynamics associated with optimization using online mirror descent and online gradient descent, we were able to validate that numerical implementations of KMD can identify conjugacies which are known to exist [24, 25, 26] (Fig. 2C). Additionally, this simple example demonstrates challenges existing approaches for comparing DNN training dynamics face, as comparing the losses and the parameter evolutions of OMD and OGD does not lead to a clear indication of the underlying equivalence (Fig. 2A, B).
@@
Leveraging our Koopman-based approach on the training dynamics of DNNs of varying architectures led to several insights. First, we found that shallow and wide fully connected neural networks have non-conjugate training dynamics (Fig. 3). This is consistent with theoretical and experimental work finding that FCN width can lead to lazy and rich training regimes [6]. This provides further evidence that our Koopman-based approach can correctly identify equivalent and non-equivalent training dynamics. In addition, we find that FCNs of intermediate width have Koopman eigenvalues that are more similar to those of wide FCNs (Fig. 3), demonstrating that our approach can provide insight beyond the wide and shallow regimes. Second, applying our framework to the dynamics of convolutional neural networks, we found transitions in the dynamics during the early phase of training, consistent with prior work [27] (Fig. 4). However, by closely examining the Koopman eigenvalues, we found non-conjugate dynamics between CNN architectures, suggesting fundamental differences in training. These distinct dynamical features are aligned with previous observations of different behaviors when training sparse CNNs [11, 38]. And third, we found that Transformers that do and do not undergo grokking have non-conjugate training dynamics (Fig. 6). By focusing on the early phase of Transformer training, we avoid trivially finding this due to differences in the training loss. Additionally, this provides evidence for the ability to detect grokking before it happens [59].
@@
Our framework is similar in spirit to an approach that categorizes iterative algorithms from a controlled dynamical systems perspective [60, 61]. However, such an approach requires access to the underlying equations to identify equivalence classes, which our Koopman-based framework does not [62, 63]. Work concurrent to ours has leveraged a similar approach to studying the dynamics of recurrent neural networks [64]. However, this work studied the dynamics of the activations and not the dynamics of training, which was the focus of this paper.
@@
Limitations. Numerical implementations that compute the KMD are only approximations to the action of the true Koopman operator. As such, they are subject to the same difficulties as other data-driven approaches. These include the selection of hyper-parameters associated with the construction of the Koopman operator, the choice of observable functions, and the number of modes considered. To mitigate the effect these limitations might have on our analysis, we used DMD-RRR, a state-of-the-art numerical approach for KMD [51], and time-delayed observables, which have been found to provide robust results across a range of domains [43, 49, 42]. Determining the existence of a topological conjugacy between two dynamical systems requires assessing whether their associated Koopman eigenvalues are sufficiently similar. While in some cases this is clear (e.g. identical Koopman eigenvalues associated with optimization using OMD and OGD – Fig. 2, distinct Koopman eigenvalues associated with training LeNet and ResNet-20 – Fig. 4), in other cases it is less apparent. To quantify these differences, we made use of the Wasserstein distance. While a natural choice, additional work remains to connect the magnitude of the Wasserstein distance to the divergence of the dynamical properties associated with training DNN models.
@@
Future directions. The ability of our Koopman-based approach to identify conjugacies between iterative optimization algorithms suggests it potential as a data-driven discovery and generation method of new classes of algorithms [65, 66, 67]. By identifying equivalent training dynamics of DNNs, it may be possible to use our approach for learning mappings that transform one DNN model to another [68]. Finally, the characterization of the Koopman eigenvalues associated with training a wide range of DNN models, varying in architecture, optimization hyper-parameters, and initialization will enable a detailed understanding of how these properties shape DNN training. Leveraging this understanding may lead to improved methods for DNN training and model development.
@@ Acknowledgments
We thank members of AIMdyn Inc., Mitchell Ostrow, Adam Eisen, Ila Fiete, and members of the Fiete Group for helpful discussion surrounding this work. This work was funded by the Air Force Office of Scientific Research (AFOSR) project FA9550-22-1-0531.
@@ A.1 Conjugacy between OMD and OGD
Let $\mathcal{K}$ be a convex set. OMD is applied to find a minimum of the function $f$ on $\mathcal{K}$, subject to some convex regularizer $R$. For each iteration of OMD, the state of the algorithm (initialized at $x(0)\in\mathcal{K})$, is updated by performing $(\nabla R)^{-1}\left(\nabla R[x(t)]-\eta\nabla f[x(t)]\right)$, where $\eta$ is the learning rate. Because this step may not be in $\mathcal{K}$, the Bregman projection operator, $\Pi_{\mathcal{K}}^{R}(x)=\text{arg min}_{y\in\mathcal{K}}D_{R}(y||x)$, is used. OGD, on the other hand, is applied to a (possibly) non-convex set $\mathcal{K}^{\prime}$ and a (possibly) non-convex function $\tilde{f}$. As with OMD, each iteration of OGD involves updating the state (initialized at $u(0)\in\mathcal{K}^{\prime}$) and projecting the update back into the set $\mathcal{K}^{\prime}$. In this case, the update is $u(t)-\eta\nabla\tilde{f}[u(t)]$, where again $\eta$ is the learning rate and the projection is done via the Euclidean projection operator, $\Pi_{\mathcal{K}^{\prime}}(x)$ (see Algorithms 1 and 2 for pseudocode implementations of both algorithms). When $\nabla\tilde{f}[u(t)]=\nabla f(q[u(t)])$ and $\mathcal{K}^{\prime}=q^{-1}(\mathcal{K})$, the outputs of the two algorithms are equivalent via the mapping $q$ (i.e., $q$ is the topological conjugacy or reparameterization). A key theorem of [25] showed that, in continuous-time, if $x(t)=q([u(t)])$, then $\frac{\partial u}{\partial t}=-\eta\nabla f(q[u(t)])$ [26].
@@
A.2 Numerical experiments

### v3

@@ 4.2 Identifying the effect of width on fully connected neural network training
Computing the Koopman eigenvalues associated with training FCNs of varying width (Fig. 3D – see Appendix C.2 for details), we find that narrow ($h=5$) and wide ($h=40$) FCNs have training dynamics that are non-conjugate, as their Koopman spectra are non-overlapping (Fig. 3E, F). This suggests that the training dynamics undergo a fundamental change as width increases. However, for FCNs with intermediate width ($h=10$), the training dynamics are more aligned with the wide FCNs (Fig. 3E, F), suggesting conjugate dynamical behavior. The dynamical difference in training narrow and wide FCNs is also supported by performing the eigenvalue shuffle analysis (Appendix B). In particular, a much larger number of the shuffles between $h=10$ and $h=40$ eigenvalues have Wasserstein distance greater than or equal to the true Wasserstein distance than between $h=5$ and $h=40$ ($81\%$ vs. $55\%$), although significance is not reached. Similar results were found when using the GeLU instead of ReLU activations [59] (Fig. S3), demonstrating that our results are consistent across FCNs with similar activation functions. Thus, we conclude that the additional improvement in performance observed when increasing the network width from $h=10$ to $h=40$ comes more from an increase in capacity, than from a change in training dynamics. Identifying this was not possible by solely comparing the loss or weights (Fig. 3A–C), demonstrating the advantage of the Koopman-based approach for identifying equivalent and non-equivalent DNN training dynamics.
To further study the behavior of FCN training dynamics, we also compared the computed Koopman spectra of $h=40$ networks trained from different random initial conditions (Appendix C.4). Prior work has proven that different random initializations of sufficiently wide FCNs converge to local minima that have no loss barrier along the linear interpolation between them, when taking into account permutation symmetry [16]. This suggests conjugate training dynamics, although this has not been explicitly shown. In support of this hypothesis, we find examples of FCNs, trained from different random initializations, with nearly identical Koopman spectra (Fig. S4A).
@@ 4.3 Identifying dynamical transitions in convolutional neural network training
Prior work has argued that CNNs undergo transitions in their training dynamics during the early part of training (i.e. the first several epochs), and that these transitions are similar across different CNN architectures [27]. However, dynamical systems based methods were not used for analysis. Instead, this observation relied on coarse-grained observables (e.g., training loss, magnitude of gradients) to define the transitions and to determine when they occur.
@@
To understand whether such results hold when considering the training dynamics at a finer-scale, we utilize our Koopman-based framework. To do this, we split the first epoch of training into windows of 100 training iterations. We compute the Koopman eigenvalues associated with dynamics that occur in each window and denote them by $\lambda_{t_{1}:t_{2}}$, where $t_{1}<t_{2}$ are the first and last training iteration in the window. We then measure the Wasserstein distance between all combinations of pairs of eigenvalues. This enables us to quantitatively assess transient dynamical behavior and identify when in the early phase of training the dynamics transition from one equivalence class to another. We apply our approach to LeNet [60], a simple CNN trained on MNIST, and ResNet-20 [61], trained on CIFAR-10 (see Appendix D.1 for details).
@@
We find that, for both LeNet and ResNet-20, the first 100 training iterations have the most distinct Koopman eigenvalues, as the Wasserstein distance between $\lambda_{0:99}$ and all other eigenvalues is large (Fig. 4A, B – bright yellow first column and row). In addition, for both LeNet and ResNet-20, the training dynamics become similar after $700$ training iterations, as the Wasserstein distance between $\lambda_{600:699}$ and $\lambda_{700:799}$ is small (Fig. 4A, B – dark blue square around diagonal in lower righthand corner). This is in agreement with the timeline found by Frankle et al. (2020) [27]. However, we additionally find that the dynamics that occur between 100 and 700 training iterations exhibit greater change for ResNet-20 than for LeNet, as there is a larger Wasserstein distance between Koopman eigenvalues. This suggests a difference in dynamical behavior between the architectures. By examining the Koopman eigenvalues associated with different training iteration windows, we find non-overlapping spectra (Fig. 4C). Performing the eigenvalue shuffle analysis (Appendix B), we find evidence that the first 4 splits of 100 training steps have statistically significant different associated Koopman eigenvalues, as the $2\%$, $2\%$, $0\%$, and $4\%$ of the shuffles had Koopman eigenvalues greater than or equal to the true Wasserstein distance. This suggests a lack of topological conjugacy between the earliest training dynamics of ResNet-20 and LeNet, despite the fact that the general timeline in transitions in dynamics is similar between the architectures.
@@
Figure 4: Koopman-based framework enables identification of transitions in dynamics during the early phase of training for LeNet and ResNet-20. (A) Log10 Wasserstein distance between Koopman eigenvalues associated with LeNet training over windows of 100 training iterations during epoch 1. (B) Same as (A), but for ResNet-20 training. (C) Koopman eigenvalues associated with the dynamics that occur during training iterations intervals 0–99, 400–499, and 600–699. Dashed line denotes the unit circle.
@@
To understand how the training dynamics change over a larger span of training time, we perform the same analysis, but computing the Koopman eigenvalues from the dynamics that occur during each epoch (Appendix D.3). We find that, at this coarser grain scale, both architectures see a similar evolution of their training dynamics. In particular, we find that the first epoch has the most distinct dynamics (Fig. S5A, B – yellow first column and row), and the subsequent epochs have dynamics that become increasingly more similar (Fig. S5A, B – increasing size of dark blue blocks centered on the diagonal).
@@
Taken together, our Koopman-based analysis supports prior decomposition of the early phase of CNN training dynamics into regimes separated by transitions that occur after a similar number of training iterations across architectures [27]. However, with a finer-scale resolution of the dynamics, we additionally find that LeNet and ResNet-20 have non-conjugate training, demonstrating that the exact training dynamics are architecture-specific.
@@ 4.4 Identifying non-conjugate training dynamics for Transformers that do and do not grok
Since the discovery that Transformers trained on algorithmic data (e.g., modular addition) undergo delayed generalization (“grokking” – Fig. 5A) [28], considerable effort has been invested to understand how this arises. One particularly influential theory is that the norm of the weights at initialization plays a large role. In particular, it was shown that single layer Transformers, initialized at weights with a sufficiently small norm, have training and test loss landscapes that are “aligned”, while the same single layer Transformers, initialized at weights with a sufficiently large norm, have training and test loss landscapes that are “mis-aligned” [62]. Constraining the norm of the weights to be small prevents grokking, with train and test accuracy increasing at similar training iterations (Fig. 5B) [62].
Figure 5: Transformers that do, and that do not undergo grokking have early training dynamics that are not conjugate. (A) Train and test loss, as a function of training steps, for a Transformer model that undergoes grokking. (B) Same as (A), but for a Transformer whose training is constrained to have a constant weight norm [62]. In this case, no grokking is observed. (C) In the first 100 training steps, little difference is seen between the test loss of Transformers with and without constrained training. Lines are mean and shaded area is $\pm$ standard deviation across 20 independently trained networks. (D) Koopman eigenvalues associated with the dynamics that occur over the first 100 training iterations for Transformers that do, and that not undergo grokking.
What role the training dynamics play in grokking remains less understood. In particular, the extent to which constraining the weight norm changes the training dynamics (which could shed additional light on grokking) has yet to be explored. We therefore compute the Koopman eigenvalues associated with the training of constrained and unconstrained Transformers on modular addition (Appendix E). We use the dynamics from the earliest part of training, namely the first 100 training iterations (Fig. 5C). We do this to avoid trivially seeing a difference, given the small weight changes that Transformers which undergo grokking make when the training accuracy is high.
@@
These results suggest a non-conjugacy in the training dynamics of Transformers that do, and those that do not undergo grokking. In particular, constraining the weight norm appears to lead to more stable training dynamics, which may be due to the selection of a better subnetwork to train [63]. Additionally, these results suggest that it may be possible to identify grokking before it happens [64].
@@ 5 Discussion
Motivated by the need for quantitative methods that can determine when the training dynamics of DNN parameter trajectories are equivalent, we utilized advances in Koopman operator theory to develop a framework that can identify topological conjugacies. This Koopman based identification of conjugate training dynamics is invariant to permutations of DNN parameters (Eq. 5), a necessary feature for methods used to study DNN training dynamics [15, 16, 17]. By applying our approach to the dynamics associated with optimization using OMD and OGD, we were able to validate that numerical implementations of KMD can identify conjugacies which are known to exist [24, 25, 26] (Fig. 2C). Additionally, this example demonstrates challenges existing approaches for comparing DNN training dynamics face, as comparing the losses and the parameter evolutions of OMD and OGD does not lead to a clear indication of the underlying equivalence (Fig. 2A, B).
@@
Leveraging our Koopman-based approach on the training dynamics of DNNs of varying architectures led to several insights. First, we found evidence that shallow and wide FCNs have non-conjugate training dynamics (Fig. 3). This is consistent with theoretical and experimental work showing that FCN width can lead to lazy and rich training regimes [6]. This provides further evidence that our Koopman-based approach can correctly identify equivalent and non-equivalent training dynamics. In addition, we find that FCNs of intermediate width have Koopman eigenvalues that are more similar to those of wide FCNs (Fig. 3), demonstrating that our approach can provide insight beyond the wide and shallow regimes. Second, applying our framework to the dynamics of CNNs, we found transitions in the dynamics during the early phase of training, consistent with prior work [27] (Fig. 4). However, by closely examining the Koopman eigenvalues, we found non-conjugate dynamics between different CNN architectures, suggesting fundamental differences in training. These distinct dynamical features are aligned with previous observations of different behaviors when training sparse CNNs [11, 38]. And third, we found that Transformers that do, and that do not undergo grokking have non-conjugate training dynamics (Fig. 5). By focusing on the early phase of Transformer training, we avoid trivially finding this due to differences in the training loss. Additionally, this provides evidence for the ability to anticipate grokking before it happens [64].
@@
Our framework is similar in spirit to an approach that categorizes iterative algorithms from a controlled dynamical systems perspective [65, 66]. However, such an approach requires access to the underlying equations to identify equivalence classes, which our data-driven, Koopman based framework does not [67, 68]. Work concurrent to ours has leveraged a similar approach to study the dynamics of recurrent neural networks [69]. However, Ostrow et al. (2023) [69] studied the dynamics of the activations and not the dynamics of network parameters, which is the focus of this paper.
@@
Limitations. Numerical implementations that compute the KMD are only approximations to the action of the true Koopman operator. As such, they are subject to the same difficulties as other data-driven approaches. These include the selection of hyper-parameters associated with the construction of the Koopman operator, the choice of observable functions, and the number of modes considered. To mitigate the effect these limitations might have on our analysis, we used DMD-RRR, a state-of-the-art numerical approach for KMD [53], and time-delayed observables, which have been found to provide robust results across a range of domains [44, 45, 51]. Determining the existence of a topological conjugacy between two dynamical systems requires assessing whether their associated Koopman eigenvalues are sufficiently similar. While in some cases this is clear (e.g. identical Koopman eigenvalues associated with optimization using OMD and OGD – Fig. 2, distinct Koopman eigenvalues associated with training LeNet and ResNet-20 – Fig. 4), in other cases it is less apparent. To quantify these differences, we made use of the Wasserstein distance and attempted to compute significance with a randomized shuffle control. While a natural choice, additional work remains to connect the magnitude of the Wasserstein distance to the divergence of the dynamical properties associated with training DNN models.
@@
Future directions. The ability of our Koopman-based approach to identify conjugacies between iterative optimization algorithms suggests its potential for data-driven discovery and generation of new classes of algorithms [70, 71, 72]. By identifying equivalent training dynamics of DNNs, it may be possible to use our approach for learning mappings that transform one DNN model to another [73]. Finally, the characterization of the Koopman eigenvalues associated with training a wide range of DNN models varying in architecture, optimization hyper-parameters, and initialization will enable a detailed understanding of how these properties shape DNN training. Leveraging this understanding may lead to improved methods for DNN training and model development.
@@ Acknowledgments
We thank members of AIMdyn Inc., Mitchell Ostrow, Adam Eisen, Ila Fiete, and members of the Fiete Group for helpful discussion surrounding this work. We thank the anonymous NeurIPS reviewers for their thorough and detailed feedback, which strengthened or work. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-22-1-0531.
@@ A.1 Conjugacy between OMD and OGD
Here we outline the conjugacy between Online Mirror Descent (OMD) and Online Gradient Descent (OGD). We follow the notation and framing presented in Ghai et al. (2022) [26].
Let $\mathcal{K}$ be a convex set. OMD is applied to find a minimum of the function $f$ on $\mathcal{K}$, subject to a convex regularizer $R$. For each iteration of OMD, the state of the algorithm (initialized at $x(0)\in\mathcal{K})$, is updated by performing $(\nabla R)^{-1}\left(\nabla R[x(t)]-\eta\nabla f[x(t)]\right)$, where $\eta$ is the learning rate. Because this step may not be in $\mathcal{K}$, the Bregman projection operator, $\Pi_{\mathcal{K}}^{R}(x)=\text{arg min}_{y\in\mathcal{K}}D_{R}(y||x)$, is used. OGD, on the other hand, is applied to a (possibly) non-convex set $\mathcal{K}^{\prime}$ and a (possibly) non-convex function $\tilde{f}$. As with OMD, each iteration of OGD involves updating the state (initialized at $u(0)\in\mathcal{K}^{\prime}$) and projecting the update back into the set $\mathcal{K}^{\prime}$. In this case, the update is $u(t)-\eta\nabla\tilde{f}[u(t)]$, where again $\eta$ is the learning rate and the projection is done via the Euclidean projection operator, $\Pi_{\mathcal{K}^{\prime}}(x)$ (see Algorithms 1 and 2 for pseudocode implementations of both algorithms). When $\nabla\tilde{f}[u(t)]=\nabla f(q[u(t)])$ and $\mathcal{K}^{\prime}=q^{-1}(\mathcal{K})$, the outputs of the two algorithms are equivalent via the mapping $q$ (i.e., $q$ is the topological conjugacy or reparameterization). A key theorem of [25] showed that, in continuous-time, if $x(t)=q([u(t)])$, then $\frac{\partial u}{\partial t}=-\eta\nabla f(q[u(t)])$ [26].
@@
A.2 Bisection method
Let $\mathcal{K}^{\prime\prime}=[c,d]^{d}$ where $f(c)<0$, $f(d)>0$, and there exists only one $z^{*}\in[c,d]^{d}$ s.t $f(z^{*})=0^{d}$. Let $a(0),b(0)\in\mathcal{K}^{\prime\prime}$ s.t. $f[a(0)]<0$ and $f[b(0)]>0$. Define $z(t)=[a(t)+b(t)]/2$. For each iteration of the Bisection Method (BM), if $f[z(t)]<0$, then $[a(t),b(t)]$ is updated to $[z(t),b(t)]$. Otherwise $[a(t),b(t)]$ is updated to $[a(t),z(t)]$ (see Algorithm 3 for pseudocode implementation).
Because not one but three variables are being updated at each iteration of the BM, [$a(t)$, $b(t)$, $z(t)$], the BM can exhibit several distinct properties from OMD and OGD. First, how much the updated $z(t+1)$ differs from $z(t)$ depends on $a(t)$ and $b(t)$. Thus, $||z(t+1)-z(t)||_{2}$ can be much larger than the steps OMD and OGD takes. This global property enables it to escape local minima that OMD/OGD get stuck in, but can also lead to increases in loss. This behavior can be seen in Fig. 2B. Second, because $f[a(t)]<0<f[b(t)]$ and because $z(t+1)$ is either $a(t)$ or $b(t)$, the outputs of the BM [i.e., $z(t)$] can flip sign. This kind of oscillatory behavior can be seen in Fig. 2A. This would not happen with OMD or OGD, assuming a sufficiently small learning rate. For these reasons, we expect that the bisection method is not conjugate to OMD or OGD. The distinct Koopman spectra (Fig. 2C) demonstrate that our framework properly identifies this non-conjugacy.
Algorithm 3 Bisection Method
Input: $f,a(0)\in\mathcal{K}^{\prime\prime},b(0)\in\mathcal{K}^{\prime\prime}$ s.t. $f[a(0)]<0<f[a(b)]$
for $t=0,...,T-1$ do
    $z(t)=[a(t)+b(t)]/2$
    if $f[z(t)]<0$ then
         $a(t+1)=z(t)$
         $b(t+1)=b(t)$
    else
         $a(t+1)=a(t)$
         $b(t+1)=z(t)$
    end if
end for
A.3 Numerical experiments